OpenStudy (anonymous):

An engineer is surveying a building with a device that is 6 feet above the ground. He is 60 feet from the building and is looking with an angle of elevation of 45 degrees. How tall is the building?

OpenStudy (anonymous):

tan 45 = h/ 60 -> h = 60 Thus the whole height 60 + 6 = 66 ft

OpenStudy (anonymous):

Thank you (: